# DriveSuprim: Towards Precise Trajectory Selection for End-to-End Planning

### v2

@@ Appendix B Visualization Results
Figure 5: Visualization results across various challenging scenarios. The first row shows the results of Hydra-MDP++, while the second row presents our method. In each example, the green trajectory represents the ground truth from the human expert, the red trajectory is generated by the classic selection-based method, and the blue trajectory is produced by DriveSuprim. Zoom in for a better view.
Appendix C Supplementary Experiment Result
C.1 Model Performance with Different FOV Settings
We show the model performance with different FOV settings in Tab 11. The FOV is revealed by the number of cameras. As shown in Fig 6, five cameras are involved in these settings: $f$ (front camera), $l_{0}$ (front-left camera), $l_{1}$ (left camera), $r_{0}$ (front-right camera), and $r_{1}$ (right camera). The 1-camera setting uses $f_{0}$ as input, the 3-camera setting uses $l_{0}$, $f_{0}$ and $r_{0}$ as input, while the 5-camera setting uses $l_{1}$, $l_{0}$, $f_{0}$, $r_{0}$ and $r_{1}$ as input. Experimental results show that the 3-camera setting leads to the best performance.
@@ C.1 Model Performance with Different FOV Settings
Table 11: Results with different FOVs (number of cameras).
1 Camera 97.5 97.8 99.4 99.8 91.2 96.7 96.9 98.3 76.2 85.4
3 Cameras 98.4 98.6 99.6 99.8 90.5 97.8 97.0 98.3 78.6 87.1
5 Cameras 98.4 98.4 99.6 99.8 90.5 97.8 97.0 98.3 78.4 86.9
Appendix D Limitations and Future Work
In this paper, we evaluate the performance of DriveSuprim on the challenging NAVSIM benchmark. Our two-stage coarse-to-fine trajectory filtering approach proves effective in enhancing model performance. However, extending this filtering strategy to a multi-stage setting does not yield additional improvements, and the inference speed remains suboptimal.
In future work, we will further investigate multi-stage refinement techniques and develop more efficient planning methods.

### v3

@@ D. Visualization
Smoothing Method EPDMS $\uparrow$
Label Smoothing 82.5
Temperature-scaled CE 82.7
Self-distillation 83.1
Table 13: Comparison of different smoothing methods.
Model Backbone EPDMS $\uparrow$ Param. $\downarrow$ FPS $\uparrow$
HydraMDP++ ResNet34 81.4 53M 32.0
HydraMDP++ V2-99 85.1 103M 24.0
HydraMDP++ Vit-L 85.6 337M 14.4
DriveSuprim ResNet34 83.1 61M 27.2
DriveSuprim V2-99 86.0 110M 22.1
DriveSuprim Vit-L 87.1 345M 12.5
Table 14: Comparison of parameter and inference speed.
1 Camera 97.1 96.4 99.4 99.5 88.6 96.2 96.0 98.3 76.2 82.9
3 Cameras 97.5 96.5 99.4 99.6 88.4 96.6 95.5 98.3 77.0 83.1
5 Cameras 97.3 96.8 99.4 99.5 88.3 96.4 96.0 98.3 76.6 83.2
Table 15: Results with different FOVs (number of cameras).
@@
Figure 6: Visualization results across various challenging or sharp turning scenarios. In each example, the green trajectory represents the ground truth from the human expert, the red trajectory is generated by the classic selection-based method, and the blue trajectory is produced by DriveSuprim. Zoom in for a better view.
Trajectory distribution comparison
Fig. 7 illustrates the trajectory frequency distribution in both the original dataset and the dataset enhanced with our rotation-based augmentation method. Using the ground-truth PDM score, we count the trajectory that either scores above 0.99 or ranks among the top-three-highest scores in each scenario. We then normalize the frequency by setting the most frequent trajectory to 1. Results show that in the original dataset, the high-score trajectories are predominantly concentrated in the forward or near-forward direction, whereas in the augmented dataset, trajectories across all directions appear with similar frequency.
E. Limitations and Future Work
In this paper, we evaluate the performance of DriveSuprim on the NAVSIM and Bench2Drive benchmark. Our two-stage coarse-to-fine trajectory filtering approach proves effective in enhancing model performance. However, extending this filtering strategy to a multi-stage setting does not yield additional improvements, and the model still has potential to improve performance under challenging or corner cases. In the future, we will further investigate multi-stage refinement techniques and explore reinforcement learning for a more robust planning system.
Acknowledgments
This work was supported by National Natural Science Foundation of China (No. 62427819) and the Science and Technology Commission of Shanghai Municipality (No. 24511103100).